OpenStudy (anonymous):

Use cylindrical coordinates to evaluate the triple integral ∫∫∫sqrt(x2+y2)dV where E is the solid bounded by the circular parabaloid z=9−16(x2+y2) and the xy plane.

OpenStudy (anonymous):

This was the integral that I set up: \[\int\limits_{0}^{2\pi}\int\limits_{0}^{3/4}\int\limits_{0}^{9-16r^2}rdzdrd \theta\] And I was not able to wind up with the correct answer.

OpenStudy (anonymous):

you do z first

OpenStudy (anonymous):

you integrate and get z b/c u only have a dz

OpenStudy (anonymous):

so u get 9-16r^{2} rdrdtheta

OpenStudy (anonymous):

I know I just think that my second limit of integration might be wrong. I know how to evaluate the integral.

OpenStudy (anonymous):

then for r you get 9r-16r^3 drdtheta

OpenStudy (anonymous):

\[\int\limits_{0}^{3/4} 9r-16r^{3} drd\]

OpenStudy (anonymous):

\[\frac{ 9r^2 }{ 2} - 4r^4 \]

OpenStudy (anonymous):

from 0 to 3/4

OpenStudy (anonymous):

no there is a function r, so it should be r^2. then int[9r2-16r4]0-3/4 *2pi

OpenStudy (anonymous):

r3-16/5r5]0-3/4 * 2pi is the final answer.